OpenStudy (lilmike234):

What is the equation of an ellipse centered at (5,-1) having a horizontal minor axis of length 4 and a major axis of length 6?

OpenStudy (anonymous):

do you know the general form of an ellipse with center \((h,k)\)?

OpenStudy (anonymous):

just asking is all if the answer is "NO" i will show you

OpenStudy (lilmike234):

No I don't

OpenStudy (anonymous):

no one likes these conic section problems but they are not that hard

OpenStudy (anonymous):

general form of ellipse with center \((h,k)\) is \[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\] you are already given \((h,k)\) all you need now is \(a\) and \(b\)

OpenStudy (lilmike234):

(x-5)^2/36+(y-1)^2/16=1 was the equation I got is this correct ?

OpenStudy (anonymous):

could be let me check

OpenStudy (anonymous):

no but you have the right idea

OpenStudy (lilmike234):

Or was is suppose to be (x+5)^2/9+(y+1)^2/4=1 ?

OpenStudy (anonymous):

first off "horizontal minor axis" is shorter than the "vertical major axis" that means it looks like this 

OpenStudy (anonymous):

second answer is even closer, you got the deominators right, but they are backwards

OpenStudy (anonymous):

half of 6 is 3 and \(3^2=9\) likewise \(2^2=4\) but since it is oriented the other way, the larger number should be under the \(y\) term not the x terms

OpenStudy (lilmike234):

So (x-5)^2/9+(y-1)^2/4=1?

OpenStudy (anonymous):

again the larger number should be under the \(y\) term since the major axis is vertical

OpenStudy (anonymous):

\[\frac{(x-5)^2}{4}+\frac{(y+1)^2}{9}=1\]